OpenStudy (anonymous):

Planes P and Q leave an airport at the same time and travel in opposite directions. Plane P travels 50 miles per hour faster than plane Q. How fast is plane P traveling if they are 1,600 miles apart after 2 hours? (A) 350 (B) 375 (C) 425 (D) 475 (E) 500

OpenStudy (gw2011):

I got and answer of 425 which is (c). Here is how I arrived at the answer: d=(r)(t) This is the distance formula P=50+Q This is plane P's travel rate d=(r)(t) 1600=(50+Q)(2)+(Q)(2) 1600=100+2Q+2Q 1600=100+4Q 1500=4Q 375=Q This is Q's rate P=50+Q P=50+375 P=425 miles per hour

OpenStudy (anonymous):

thx

OpenStudy (gw2011):

You're welcome